ENERGY SPLITTING PATTERNS AND STATISTICAL WEIGHTS FOR THE BENZENE DIMER

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; UNDINE ERLEKAM, GERARD MEIJER, MELANIE SCHNELL and GERT VON HELDEN, Fritz-Haber-Institut der Max-Planck-Gesellschaft, Faradayweg 4-6, D-14195 Berlin, Germany.

Based on recent ab initio results,^{a} the benzene dimer is predicted to have a distorted T -shaped equilibrium structure; one benzene ring forming the “stem” of the “T ” and the other forming the “cap.” The barriers for several tunneling pathways, such as stem-bending, cap-torsion and stem-torsion, have been calculated. In this talk the energy splitting patterns, and nuclear spin statistical weights, that characterize such tunneling pathways will be discussed. Comparison with high resolution spectra will be made for (C_{6}H_{6})_{2}, (C_{6}D_{6})_{2}, and C_{6}H_{6}-C_{6}D_{6}.

aR. Podeszwa, R. Bukowski, and K. Szalewicz J. Phys. Chem. A 110, 10345, (2006) and R. A. DiStasio Jr., G. von Helden, R. P. Steele, and M. Head-Gordon Chem. Phys. Lett. (2007), doi: 10.1016/j.cplett.2007.02.034.